Sir, an insect of mass 'm' is initially at rest at one end of a stick of length 'L' and mass 'M' , which rests on a smooth horizontal floor.The cofficient of friction between the insect and stick is 'k'. The insect starts moving to reach the other end in shortest possible time. Q1.Find the time taken to reach the other end? Q2.What is the magnitude of linear momntum of insect at time "t"? Q3.What is the maximum speed aquired by the stick?? please sir while answering do explain ur approach..!!

Sir,


an insect of mass 'm' is initially at rest at one end of a stick of length 'L' and mass 'M' , which rests on a smooth horizontal floor.The cofficient of friction between the insect and stick is 'k'. The insect starts moving to reach the other end in shortest possible time.


ย 


Q1.Find the time taken to reach the other end?


Q2.What is the magnitude of linear momntum of insect at time "t"?


Q3.What is the maximum speed aquired by the stick??


ย 


please sir while answering do explain ur approach..!!

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Ananya,

friction force is kmg

accelearation of insect is kg

accelearation of rod is kmg/M

acceleration of insect wrt rod is kg+kmg/M=kg(1+m/M)=a

time taken by the insect to re3ach the other end is t=√(2*L/a)

linear momentum of insect at time t = mv =m*a*t

maximum speed acquired by the rod= u

momemtum conservation: Mu=mv=m*√(2*L*a)

u=(m/M)*√(2*L*a). where a=kg(1+m/M).